# OptiMUS-0.3

### v1

@@ 4 Methodology
This section details the design of OptiMUS-0.3. The workflow of OptiMUS-0.3 is outlined in Algorithm 1. For brevity, the term clause refers to a constraint or objective within the problem.
OptiMUS-0.3 begins by extracting parameters from the problem description (line 2). It then extracts (line 4), and formulates (line 7) the clauses. During clause formulation, the variables, their auxiliary constraints, and a connection graph (a graph that records which variables and parameters appear in each constraint) are also generated. OptiMUS uses this connection graph to ensure consistency of the formulations. The connection graph is key to performance and scalability of OptiMUS, as it allows the LLM to focus only on the relevant context for each prompt, generating more stable results. An example of the constraint formulation step appears in Fig. 4
Figure 4: The formulation process for a single constraint. OptiMUS-0.3 identifies any parameters and variables appearing in the constraint, including new variables that it may need to define. It defines new variables as needed, updates the connection graph which records which constraints use which parameters and which variables, and annotates the constraint with a formulation. (dashed lines represent new connections and variables)
Subsequently, OptiMUS-0.3 generates code snippets for all parameters, clauses, and variables (line 11), assembles these snippets into a single runnable code file (line 12), and iteratively executes and debugs to achieve the desired results (lines 13-15) or until timeout. Error correction (EC) after the extraction and formulation stages catch common errors and improve the solve rate. All prompts used in OptiMUS-0.3 appear in Concluding the reduction.
Algorithm 1 Workflow of OptiMUS-0.3
1:Input: Natural language description of problem $\mathcal{P}$
[…]
3:$\textsc{CorrectErrors}(\textup{params},Q_{params})$
[…]
5:$\textsc{CorrectErrors}(\textup{cons},Q^{(nl)}_{cons})$
6:$\textsc{CorrectErrors}(\textup{obj},Q^{(nl)}_{obj})$
[…]
8:$\textsc{CorrectErrors}(\textup{cons},Q^{(form)}_{cons})$
9:$\textsc{CorrectErrors}(\textup{obj},Q^{(form)}_{obj})$
10:$\textsc{CorrectErrors}(\textup{vars},Q_{vars})$
[…]
13:while Execute(code) throws error e do
[…]
15:end while
16:res $\leftarrow$ Execute(code)
Algorithm 2 CorrectErrors(X, Q)
1:Input: a set of JSON objects $X$, and a list $Q$ of tuples of the form $(q_{llm},q_{user},f)$ where $q_{llm}$ is an LLM prompt template, $q_{user}$ is a user prompt template, and $f$ is a callback function.
2:for $x\in X$ do
3:     for $(q_{llm},q_{user},f)\in Q$ do
4:         $\mathrm{prompt}\leftarrow q_{llm}.\mathrm{format}(x)$
5:         $\mathrm{response},\mathrm{confidence}\leftarrow\mathrm{LLM}(prompt)$
6:         if $\mathrm{confidene}\leq\lambda$ then $\triangleright$ If the LLM is unsure what to do, then ask the user
7:              $\mathrm{user\_prompt}\leftarrow q_{user}.\mathrm{format}(\mathrm{response},x)$
8:              $\mathrm{response}\leftarrow\mathrm{AskUser}(\mathrm{user\_prompt})$
9:         end if
10:         $x\leftarrow f(\mathrm{response},x)$triangleright Update the object
11:     end for
12:end for
13:return $X$
@@ 4.1 State
  • Parameters: OptiMUS-0.3 can choose a symbol for each parameter, infer its shape, and define the parameter if it is not explicitly included in the problem statement. Numerical data from the problem statement is omitted from the parameter list and stored separately for later use, ensuring that the parameter list remains concise and easy to include in future prompts.
  • Clauses: Each clause (objective or constraint) contains a natural language description, a formulation, and a definition code.
@@
  • Background: The background is a short string that explains the real-world context of the problem. Including this string in every prompt helps improve common sense reasoning.
@@
The state is initialized as an empty object, and is completed step by step throughout the process. An example of a completed state for a factory optimization problem appears in Fig. 5.
@@ 4.2 Error Correction
To build a trustworthy and reliable system using LLMs, it is important to mitigate the impact of LLM hallucinations. In the context of optimization modeling, an LLM might confidently generate incorrect parameters, redundant mathematical constraints, or erroneous code. To address this issue, OptiMUS-0.3 uses two main error correction techniques: reflective prompts and confidence-based user feedback.
Figure 6: OptiMUS-0.3 can fix its parameter and variable identification errors when prompted “Is the value of $P$ known or not?”
@@ 4.2.1 Reflective prompts
LLMs can often identify and fix their mistakes by using reflective prompts Shinn et al. (2023). We analyzed the most common types of mistakes made by LLMs at each step of the process and designed reflective prompts to address each one. Using these prompts significantly reduces the modeling error rate. The resulting substantial improvement to OptiMUS substantiates the claim that domain-specific knowledge can be used to develop LLM applications that outperform general-purpose LLMs (see Table 3). The reflective prompts used by OptiMUS-0.3 are as follows:
Parameter Extraction.
The model often confuses parameters with variables, misidentifies parameter shapes or misses some parameters. We use the following reflective prompt to correct these errors:
Figure 7: OptiMUS-0.3 can fix parameter identification errors when prompted “Is the value of $P$ known or not?”
  • “Is the value of $P$ known or not? Based on that, is it a parameter or a variable?” (see Fig. 7)
Constraint Extraction.
During constraint extraction, the model sometimes extracts trivial or vague constraints and may regenerate a previously generated constraint in different language. We use the following reflective prompts to correct these errors:
Does constraint $C$ need to be explicitly modeled in the mathematical formulation? (see Fig. 6) Here’s the list of all constraints: $C_{1},C_{2},\ldots,C_{k}$. Are any of them redundant?
@@
Clause Modeling. Once the clauses are extracted, OptiMUS-0.3 models them, defining variables and auxiliary constraints if needed. Often, the model generates incorrect mathematical formulations for constraints. We use the following reflective prompts to detect and correct these errors:
Are units the same for both sides of $C$? (see Fig. 8) Is $C$ a correct constraint in the context of MILP solvers? (see Fig. 9) Does constraint $C$ involve any variables, or is it only defined over parameters?
Figure 9: OptiMUS-0.3 can fix its constraint modeling errors if by asking Is constraint $C$ a correct constraint in the context of MILP solvers?
Figure 10: OptiMUS-0.3 can fix its parameter and variable identification errors when prompted “Is the value of $P$ known or not?”
OptiMUS-0.3 uses similar reflective prompts to perform error checking for the optimization objective. For variables, we use the following prompt:
Is the value of $V$ known or not? Based on that, is it a parameter or a variable? (see Fig. 10)
@@ 4.2.2 Confidence-based user/LLM feedback
In the context of real-world optimization modeling, it is exceeding rare for a natural-language problem description to correspond to an unambiguous MILP formulation. Rather, the process of optimization modeling generally involves many cycles of formulation and user feedback to arrive at a correct model for the problem. OptiMUS-0.3 simulates this process in order to detect errors and improve its solve rate. OptiMUS-0.3 uses a confidence-based reflective method to decide when to request help with its formulation, either from a user (if available) or another (more powerful) LLM; for simplicity, we refer to this partner as the user below.
First, OptiMUS-0.3 scores its confidence in its output:
How confident are you in your response (from 1 to 5)?
If the confidence score is less than 4, OptiMUS-0.3 asks the user to help. The user 11 can decide to either keep the constraint, remove it, or modify it.
@@
Figure 11: Asking GPT-4o to provide feedback when LLama-3-70B’s confidence is low in identifying a constraint.
@@ 5 Experiments
In this section, we conduct a comprehensive evaluation of OptiMUS-0.3. We begin by detailing the datasets used in our experiments and showcase the superior performance of OptiMUS-0.3 across these datasets, highlighting its strengths. An ablation study demonstrates the impact of different system components on our results, and a sensitivity analysis probes the internal dynamics of OptiMUS-0.3. We conclude this section by identifying failure cases and potential areas for further improvement.
@@ 5.1 Overall Performance
To evaluate the overall performance of OptiMUS, we compare it with standard prompting, Reflexion, and Chain-of-Experts (CoE) Shinn et al. (2023), Xiao et al. (2023). Reflexion is the highest-performing general-purpose framework and CoE is the state-of-the-art method for natural-language optimization modeling. Three main metrics have been used in the literature: accuracy, compilation error (CE) rate, and runtime error (RE) rate. However, a method can generate a totally irrelevant short code that runs, or fix runtime and complication errors by completely removing relevant sections of the code. Hence, we only compare the models’ accuracy. Accuracy is defined as the number of instances correctly solved (An instance is considered as correctly solved only if the code runs successfully and the optimal value is correct. Optimal values are obtained from the dataset or by solving the problems manually). Results are presented in Table 3. OptiMUS-0.3 outperforms all other methods in all datasets by a large margin. This remarkable performance improvement highlights the importance of modularity and structure compared to a single prompt to solve complex problems using LLMs.

### v2

@@ 4 Methodology
An optimization problem proposes to maximize or minimize a given function $f$ over a set of allowable inputs ${\cal X}$. This paper focuses on modeling mixed-integer linear optimization problems (MILP) where $f$ is a linear function and the set of allowable inputs ${\cal X}$ can be represented by a set of linear inequalities over integer and real valued variables. A given MILP comprises three key components: (1) a set of $n$ discrete and real variables (i.e., the inputs that can be changed), (2) a set of linear inequalities and equalities that define different constraints on the input variables, and (3) a linear objective function that defines what the problem aims to minimize or maximize. A MILP can be written mathematically as
$\underset{\{x\}}{\text{minimize}}\quad$ $\sum_{j=1}^{n}c_{j}x_{j}$
subject to $\sum_{j=1}^{n}a_{ij}x_{j}\leq b_{i},\quad i=1,\ldots m_{1}$
$\sum_{j=1}^{n}a_{ij}x_{j}=b_{i},\quad i=1,\ldots m_{2}$
$x\in\mathbb{Z}^{n-r}\times\mathbb{R}^{r}.$
A feasible point $x^{*}$ that minimizes the objective function is called a (optimal) solution and has associated optimal objective value $c^{T}x^{*}$. For brevity, we refer to the equations defining constraints and the objective as different clauses in the optimization model. A parameter of an optimization model is any numeric coefficient in the optimization model (i.e., elements of $A$ or $c$). The goal of the OptiMUS system is to go from a natural language description of an optimization problem (e.g., a paragraph of text describing a business problem), which we call the problem description, to a decision (assigment of numerical values to each variable) that solves the problem.
OptiMUS-0.3 decomposes these two tasks into a sequence of LLM-powered steps that incrementally constructs an optimization model and implements it. An overview of the complete workflow of OptiMUS-0.3 is outlined in Fig. 1. At a high level, OptiMUS-0.3 begins by 1) identifying parameters from the problem description and writing a natural-language description of each clause in the problem. It then 2) formulates both the clauses and decision variables of the optimization model with mathematical precision, using  code. Finally, 3) OptiMUS-0.3 uses the  model of the optimization problem and the extracted parameters to generate code snippets for each clause in python and assembles the snippets into a single runnable code file. By running this file, OptiMUS-0.3 outputs an optimal solution and optimal objective value. Throughout this process, OptiMUS-0.3 maintains a state that documents what is known about the optimization model, as well as a connection graph that tracks which variables appear in which constraint. Each LLM-powered component of OptiMUS-0.3 also incorporates an Error Checking (EC) module that catches common errors and improves the accuracy of the modeling and implementation.
Our current implementation writes code that uses Gurobi (Achterberg 2019), a leading MILP solver, and its associated python API gurobipy. However, we have designed OptiMUS-0.3 to be easy to extend to other modeling languages and backends. We expect it to perform equally well with other solvers (e.g., CPLEX, SCIP) that have well-documented APIs, as well as with other modeling languages such as cvxpy (Diamond and Boyd 2016).
Many of the advances in optimization solvers in the last several decades rely on the fact that solvers are faster when they can exploit particular structures within the optimization problem. One of the main tasks of an optimization expert is to identify these structures so as to choose an appropriate solver or parameter settings. However, detecting useful structures in a given optimization problem can be very challenging, since it requires both a deep understanding of the mathematical structure of the problem and the capabilities of existing solvers. In view of this challenge, OptiMUS-0.3 includes two specialized optimization modules: 1) a Structure Detection Agent, that detects variables (e.g., Indicator variables) and constraints (e.g., SOS constraints) with special structure that can be exploited by the solver, and 2) a customized optimization coding agent, that can leverage callbacks in the solver and perform basic column and constraint sifting to help improve the runtime of the code. These elements represent a first step towards automating more advanced decomposition algorithms with an automated modeling framework.
The remainder of this section details the key components of the OptiMUS-0.3 framework. Section 4.1 describes the underlying state and connection graph of the system. Section 4.2 describes a sample LLM component of the system. The remaining sections describe specialized modules that improve the performance of the system, including error correction and debugging (Section 4.3), specialized modeling tools (Section 4.4), and advanced optimization coding (Section 4.5). The components described in this Section have been implemented and released as open-source code111included in the supplementary material and are publicly available to use through a web application, described in Section 6.
Figure 1: Optimus uses a sequential process with error correction (green circles) to model and solve optimization problems. It extracts the parameters and clauses, formulates them, generates the code for each, and finally synthesizes complete code and runs it. If the code raises an error, Optimus uses an iterative debugging loop to fix the code. Red boxes provide sample outputs of each step of the pipeline. Blue boxes indicate LLM-based components, and yellow boxes indicate deterministic components.
@@ 4.1 State
  • Parameters: OptiMUS-0.3 can choose a symbol for each parameter, infer its shape, and define the parameter if it is not explicitly included in the problem statement. Numerical data from the problem statement is replaced by a symbol and stored separately for later use, ensuring that the parameter list remains concise and easy to include in future prompts.
  • Clauses: Each clause (objective or constraint) consists of a natural language description, a formulation, and code implementing the clause.
@@
  • Background: The background is a short string that explains the real-world context for the problem. Including this string in every prompt helps improve common sense reasoning.
  • Connection Graph: The connection graph is a bipartite graph $G=(V,E)$ that links clauses to their associated parameters and variables. Formally, for every variable $x_{j}$ (clause $c_{i}$) in our formulation we create a node $v_{x_{j}}$ ($v_{c_{i}}$). In total, the connection graph consists of $n+m+1$ nodes $v\in V$ corresponding to $n$ variables, $m$ constraints and the objective. We create an edge $e\in{E}$ between a variable node $v_{x_{j}}$ and a clause node $v_{c_{i}}$ if the variable participates in the clause (i.e., $(v_{x_{j}},v_{c_{i}})\in{E}$ if $a_{ij}\neq 0$). We build the connection graph iteratively by adding nodes and edges as more constraints and variables are formulated. This structure is utilized during the coding phase to ensure that only the relevant parameters and constraints are passed to the LLM.
@@
The state is initialized as an empty object, and is completed step by step throughout the process. An example of the completed state for a factory optimization problem appears in Fig. 2.
4.2 Sample LLM Component
OptiMUS-0.3 uses LLMs as a flexible tool to perform various tasks for the overall system, including extracting parameters, modeling clauses, and coding the components (see Figure 1). In the interest of brevity, this section outlines the design of a typical LLM component with the OptiMUS framework. Each LLM component is governed by a natural language directive called a prompt. Within OptiMUS-0.3, each prompt includes the following key pieces of information:
  • Task Description: We describe the specific task, e.g., “Your job is to extract natural language constraints for this paragraph defining an optimization problem”.
  • Problem Context: We include relevant information about the problem and the current solution in the prompt. During the formulation step, this involves the specific clause being addressed, along with all parameters and variables defined so far. During the coding step, this involves the clause and its formulation, supplemented by related parameters and variables which are dynamically loaded using the connection graph.
  • Examples: We include a fixed set of sample outputs for the task in the prompt to help the LLM perform In-Context Learning (Dong et al. 2024). For modules that output python code, we also provide examples detailing the functionality of the gurobipy API.
For a comprehensive list of prompts used in OptiMUS-0.3 see our code222included in supplementary materials. We also experimented with incorporating Retrieval Augmented Generation (RAG) into OptiMUS. RAG chooses in-context learning examples that are considered most likely to be helpful for a given challenge, often by choosing examples that minimize some distance metric to the input task. However, RAG had inconsistent impact on performance (see Appendix 10) and is not used in the final version of OptiMUS-0.3.
@@ 4.3 Error Correction
To build a trustworthy and reliable system using LLMs, it is important to mitigate the impact of LLM hallucinations. In the context of optimization modeling, an LLM might generate incorrect parameters, redundant mathematical constraints, or erroneous code. To address hallucinations, OptiMUS-0.3 uses two main error correction techniques: reflective prompts and confidence-based user feedback.
@@ 4.3.1 Reflective prompts
LLMs can often identify and fix their mistakes by using reflective prompts (Shinn et al. 2023), a strategy by which an LLM is asked to reflect upon its output and decide if it needs to be corrected. We analyzed the most common types of mistakes made by LLMs on the development dataset at each step of the process and designed reflective prompts to address each one (see Appendix 9 for a full list of reflextive prompts used in OptiMUS-0.3). Using these prompts significantly reduces the modeling error rate. For instance, Figure 3 shows an example where the reflexive prompt Are units the same for both sides of this constraint? can help correct a modeling error. The resulting improvement to OptiMUS is substantial (see Table 3), and substantiates the claim that domain-specific knowledge can be used to develop LLM applications that outperform general-purpose LLMs.
@@ 4.3.2 Confidence-based user/LLM feedback
In the context of real-world optimization modeling, it is exceeding rare for a natural-language problem description to correspond to an unambiguous MILP formulation. Rather, the process of optimization modeling generally involves many cycles of formulation and user feedback to arrive at a correct model for the problem. OptiMUS-0.3 simulates this process in order to detect errors and improve its solve rate. OptiMUS-0.3 can use a confidence-based reflective method to decide when to request help with its formulation, either from a user (if available) or another (more powerful) LLM.
First, following the methodology of existing work on confidence calibration in LLMs (Virk et al. 2024), we prompt the system to asses its confidence in its output on a scale of 1 to 5. If the confidence score is less than 4, OptiMUS-0.3 asks the user (or a stronger LLM) to help. The user can decide to either keep the output (e.g., constraint, code), remove it, or modify it. In the OptiMUS webapp (Section 6), this feature is used to highlight outputs that merit particular attention. In our experiments (Section 5), this feature is turned off by default to facilitate streamlined and reproducible testing without a user in the loop. See Fig. 4 for an example of how a weaker LLM (Llama-3-70B) can call on a stronger LLM (GPT-4o) when it is not confident in its solution.
@@
Figure 4: GPT-4o can provide feedback when LLama-3-70B lacks the confidence to identify a constraint.
4.3.3 Debug Code
After the complete code is synthesized and executed by OptiMUS, any errors encountered during execution are passed to the LLM for debugging. At this stage, the LLM has the ability to modify any part of the code as needed. This step is particularly useful for resolving errors that arise from inconsistencies across individual code snippets, such as slight variations in variable names or formatting. The debugging process is repeated iteratively until the code executes successfully, with a maximum of five attempts.
4.4 Structure Detector Agent
OptiMUS maintains a pool of optimization structures commonly used in optimization software. These structures can be specified explicitly in many modern optimization solver interfaces (Achterberg 2019, CPLEX User’s Manual 1987, Gamrath et al. 2016), which can enhance problem-solving performance and simplify the code. Typical examples of structure include Special Ordered Set (SOS) (Beale and Forrest 1976), indicator variables, semi-continuous variables, and piecewise-linear constraints. At least one of these structures appear in approximately 10% of the NLP4LP dataset (see Appendix 11). We expect these structures to be even more common, and thus important to exploit, in industrial applications.
To detect structure in a constraint (or variable), OptiMUS iterates through these structures and formats them into a structure detection prompt. Within each prompt, the LLM is provided with the structure description, explained by an example illustrating how the structure should be exploited. The LLM is asked to decide whether the structure is relevant to the existing formulation. Upon identifying the appropriate structure, the formulation is adjusted to highlight the problem structure. For instance, a set of constraints that indicate only one of a group of decision variables can be non-zero may be reformulated as type-1 SOS constraint. This structure is then conveyed to gurobi via the python interfaces for special structures such as SOS constraints or indicator variables. This information can either be exploited within the branch and bound solver (e.g., using SOS constraints for branching rules) or allows the solver to automatically reformulate the structure into linear constraints (i.e., a big-M formulation for indicator constraints) (Gurobi Optimization 2023). While it may seem counter-intuitive to identify structure in a MILP only to have that structure re-formulated into linear constraints, in practice we found that automated methods within Gurobi can generate tighter formulations than OptiMUS alone (e.g., a better selection for big-M values). Figure 5 shows two instances where identifying SOS constraints and indicator variables leads to faster performance than a naive implementation. For full details of these examples are included in Appendix 11. The structure detection agent is run during the formulate clauses stage of the algorithm.
Figure 5: Impact of Structure Detection on OptiMUS Performance. (Left) Speedup of solving a facility location optimization problem with indicator variables with the structure identified to the solver (indicator) versus with OptiMUS modeling of indicator variable. (Right) Speedup of solving a flight assignment problem with SOS-constraints with the structure identified to the solver (SOS) versus with OptiMUS modeling of the SOS constraints. Full details of problem instances are included in Appendix 11.
OptiMUS also maintains a problem structure pool that enumerates combinatorial optimization problems with special structure that can be exploited in fast custom solvers (e.g., SAT, network flow, routing problems). For instance, the traveling salesman problem (TSP) can be solved much more efficiently using a specialized solver such as Concorde as opposed to general purpose MILP solvers (Cook et al. 2011). OptiMUS iterates through this pool of special problem structures given an initial natural language description of the problem. When the optimization problem falls into one of the problem types in the structure pool, OptiMUS identifies the particular problem structure and suggests using a customized solver via a notification in the OptiMUS webapp.
4.5 Advanced Optimization Coding Agent
For extremely large-scale optimization problems, optimization solvers are often embedded into a high-level optimization framework and called as a subroutine: examples include decomposition algorithms such as column generation, Bender’s decomposition, and cutting plane methods. To leverage variable and constraint structure in this large-scale context, the subproblem solver must invoke an advanced solver interface (for example, call-backs, model attribute query and analysis). While we leave the problem of developing a fully automated system that can perform advanced decomposition algorithms to future work, we have added several features to OptiMUS that advance this vision. OptiMUS implements a dedicated coding agent that exploits advanced solver functionality such as callbacks, and can iteratively call a solver within a simple variable or constraint sifting scheme (also known as delayed constraint generation). Neither of these modules require explicitly reformulating the problem (e.g., generating a pricing problem within column generation), but allow for better computational performance than a naive implementation.
Similar to the structure detector agent, the optimization coding agent maintains a series of template prompts for this functionality. Within each prompt, the LLM is provided with the purpose of the template. We illustrate this functionality with an application to sifting, a simple version of constraint or column generation, in Section 4.5.1. The advanced coding agent runs during the Code Clauses stage of OptiMUS (see Fig. 1).
4.5.1 Sifting
Sifting is a large-scale linear optimization framework initially proposed in (Bixby et al. 1992). Consider the primal linear optimization problem
$\min_{x}\quad c^{\top}x\quad\text{subject to}\quad Ax=b,\quad x\geq 0,$
where $A\in\mathbb{R}^{m\times n}$ and $n\gg m$. On the primal side, sifting combines the idea of primal simplex and column generation by solving a restricted master problem,
$\min_{x_{S}}\quad c_{S}^{\top}x_{S}\quad\text{subject to}\quad A_{S}x_{S}=b,\quad x_{S}\geq 0,$
for a subset of columns $S\subseteq\{1,\ldots,n\}$. The restricted problem can be solved by any linear programming subroutine. Its dual variable $y_{S}$ will be used to price out dual infeasible columns
$I_{S}:=\{j:c_{j}-A_{j}^{\top}y<0\}.$
Variable sifting works by iteratively updating $S\leftarrow S\cup I_{S}$ untill $S=\varnothing$. The idea can be generalized to the dual side when $m\gg n$, where a restricted problem contains a subset of rows. We provide an example of constraint sifting in OptiMUS in Fig. 6 on a security-constrained unit commitment problem. For details of the problem see Appendix 12. Most of the constraints in the instance are inactive, making it a good candidate for sifting.
Figure 6: Left: constraint sifting prompt. Right: performance plot of solving to 5% gap.
@@ 5 Experiments
In this section, we conduct a comprehensive evaluation of OptiMUS-0.3. We showcase the superior performance of OptiMUS-0.3 on the NLP4LP dataset [^3] In the tests, we assume that the parameter names are given and identified in the problem description., highlighting its strengths and weaknesses. An ablation study demonstrates the impact of different system components on our results, and a sensitivity analysis probes the internal dynamics of OptiMUS-0.3. We conclude this section by identifying failure cases and potential areas for further improvement.
@@ 5.1 Overall Performance
To evaluate the overall performance of OptiMUS, we compare it with standard prompting, Reflexion, and Chain-of-Experts (CoE) (Shinn et al. 2023, Xiao et al. 2023). Reflexion is the highest-performing general-purpose framework and CoE is the state-of-the-art method for natural-language optimization modeling. We also include results from the intial conference publication of this work, which we denote as OptiMUS-0.2. Unless otherwise stated, we use GPT-4o (OpenAI 2023) as the LLM for OptiMUS and all baseline prompting methods.
Three main metrics have been used to assess the accuracy of LP modeling tools in the literature: accuracy, compilation error (CE) rate, and runtime error (RE) rate. However, a method can generate irrelevant short code that runs, or fix runtime and compilation errors by completely removing relevant sections of the code. Hence, we only compare the models’ accuracy. Accuracy is defined as the number of instances correctly solved. An instance is considered as correctly solved only if 1) the code runs successfully; 2) the optimal value is correct; 3) the optimal solution is correct. Optimal values and solutions are obtained from the dataset or by solving the problems manually. Since the output of the code does not necessarily match the format of the optimal solution in the dataset, we use an LLM to determine whether the solution generated by OptiMUS matches the optimal solution in the dataset.
Results are presented in Table 3. For both easy and hard problems, OptiMUS-0.3 outperforms all other methods in all datasets by a large margin, beating the standard GPT-4o baseline by over 40% for hard problems. This remarkable performance improvement highlights the importance of modularity and structure compared to a single prompt to solve complex problems using LLMs. OptiMUS relies on structured code outputs (OpenAI 2024, LangChain 2024), which are currently incompatible with inference-time reasoning models, such as OpenAI’s o1-preview. As of this writing, this limitation prevented us from testing these models in our experiments.